HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

THE ROTATIONAL, RO-VIBRATIONAL EXPERIMENTAL AND COMPUTATIONAL INVESTIGATION OF MONODEUTERATED CHLOROMETHANE

\underline{F. TAMASSIA}, E. CANÈ, Dipartimento di Chimica Industriale “Toso Montanari”, Università di Bologna, Viale del Risorgimento 4, 40136 Bologna (Italy); A. DE LORENZI, A. PIETROPOLLI CHARMET, P. STOPPA,

Dipartimento di Scienze Molecolari e Nanosistemi, Università Cà Foscari Venezia,

Via Torino 155, 30172 Mestre Venezia (Italy); A. ACHILLI, L. DORE, M. MELOSSO, C. PUZZARINI, Dipartimento di Chimica “Giacomo Ciamician”,

via F. Selmi 2, Università di Bologna, 40126 Bologna (Italy)

The comprehensive description of our spectroscopic analyses and ab initio calculations for the mono-deuterated chloromethane, CH_{2}D^{37}Cl, is presented. It includes the detection of ν_{4} and ν_{8} ro-vibrational fundamental bands between 1170 and 1370 cm^{−1} in the recently recorded high-resolution FTIR spectra ( P. Stoppa et al., High resolution FTIR spectrum of CH_{2}D^{37}Cl: ν_{4} and ν_{8} fundamental bands, accepted in JQSRT) and our detailed previous analysis of ν_{5}, ν_{6} and ν_{9} fundamentals between 650 and 1100 cm^{−1} (doi:10.1016/j.jqsrt.2021.107719, P. Stoppa et al., High resolution FTIR study of the ν_{5}, ν_{6} and ν_{9} fundamental bands of CH_{2}D^{37}Cl, JQSRT 270, 107719, 2021). The vibration analyses from the medium resolution (up to 0.1 cm^{−1} ) gas-phase infrared spectra of CH_{2}DCl, investigated in the region 600-9000 cm^{−1}, are also reported (doi:10.1364/AO.39.003984, A. Pietropolli Charmet, Computational, rotational and ro-vibrational experimental investigation of monodeuterated chloromethane, JQSRT 305, 108624, 2023). All most important spectral features have been assigned in terms of fundamental, overtone and combination bands, achieving an accurate description of the vibrational structure.

The rotational spectra of the ^{13}C species have been recorded for the first time in the 250-300 GHz frequency range and, for the two isotopologues ^{12}CH_{2}D^{35}Cl and ^{12}CH_{2}D^{37}Cl, the knowledge has been extended up to 520 GHz. In addition, very accurate ab initio calculations have been performed in order to derive reliable predictions of the equilibrium structure, rotational constants, and centrifugal distortion terms for the^{12/13}C and^{35/37}Cl isotopologues. The accurate semi-experimental equilibrium structure of chloromethane has been derived. This extended experimental and computational spectroscopic work will be useful to assist the detection of CH_{2}DCl in remote environments.

p-number: p123  Submitted on Thu, 20 Jul 2023 11:20:45 +0200